given that. f: 3x^2-4x+1 g: 2x-1 calculate fg(10)
Answer by funmath(2933)

given that.
f: 3x^2-4x+1
g: 2x-1
calculate fg(10)
This means evaluate the function of f times the function of g when x=10.
:
fg(x)=(3x^2-4x+1)(2x-1)
fg(10)=(3(10)^2-4(10)+1)(2(10)-1)
fg(10)=(3(100)-40+1)(20-1)
fg(10)=(300-40+1)(19)
fg(10)=(261)(19)
fg(10)=4959
:
You could have also just found f(10) and g(10) and multiplied them if you're intimidated by the size of the equation.
f(10)=261
g(10)=19
fg(10)=261*19=4959
Happy Calculating!!!